ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

20. Special Activities of the ICM 1998

1991 MS Classification: 65, 68, 73, 76 Wieners, Christian, Bastian, P., Birken, K., Johannsen. K., Lang, S., Neuss, N., Rentz-Reichert, H., Wittum, G., Universität Stuttgart, Germany UG – A flexible software toolbox for solving partial differential equations We will give an overview of the features and the capabilities of UG. Our goal is the development of a parallel software-platform for solving partial differential equation problems. State-of-the-art numerical methods have been developed and implemented for the efficient and comfortable solution of real-world application problems. Emphasis is laid on the following topics: distributed unstructured grids, adaptive grid refinement, derefinement/coarsening, robust parallel multigrid methods, various FE and FV discretizations, dynamic load balancing, mapping and grid partitioning. Some important application examples will be presented including structural mechanics, two-phase flow in porous media, Navier-Stokes problems (CFD) and density-driven groundwater flow. Reference: P. Bastian, K. Birken, K. Johannsen, S. Lang, N. Neuss, H. Rentz-Reichert, and C. Wieners, UG - a flexible software toolbox for solving partial differential equations, Computing and Visualization in Science 1 (1997), pp. 27–40.